OpenStudy (anonymous):

evaluate the limit of x-8/ x^2-8 as x->8. I got 16.

OpenStudy (anonymous):

\[\frac{ 8-8 }{ 64-8 }=\frac{ 0 }{ 56 }\]

OpenStudy (anonymous):

that's what I got I dont know how you got 16

OpenStudy (anonymous):

\[\lim_{x\to 8} \frac{x-8}{x^2 - 8}\] Is this what you meant? then i must agree with recon14193 as x approaches 8 the denominator approaches 56, but the numerator approaches 0. The fraction tends to a finite value, which is zero.